The sides of a triangle have lengths 4,8,and 10. What is the length of the shortest side of a similar triangle that has a perimeter of 33?
Answer by stanbon(75887)

The sides of a triangle have lengths 4,8,and 10. What is the length of the shortest side of a similar triangle that has a perimeter of 33?
----
Perimeter of the given triangle: 4+8+10 = 22
---
Ratio of new triangle perimeter to given triangle perimeter: 33/22
------
short side of a similar triangle = (33/22)*4 = (3/2)4 = 6
------------------
Cheers,
Stan H.
===============